The stock of Hybrid Motors, Inc. has a beta of 0.8. The market risk premium is 10 percent and the risk free rate is 6 percent. What is the expected return on this stock?
Answer #1

Beta of the stock = β = 0.8

Market Risk premium = RM - RF = 10%

Risk-free rate = RF = 6%

Expected return on the stock is calculated using the CAPM equation

Expected return on stock = E[R] = RF + β*(RM - RF) = 6% + 0.8*10% = 14%

Answer -> Expected return on the stock = 14%